A vessel containing water is kept on a weighing machine. The
machine registers a weight of 100 N. A wooden block having cubical
shape and edge length 10 cm is gently put in water. Water does not
spill out of the vessel. Now it observed that the weighing machine
registers a weight of 106 N. Calculate the depth to which the block
will submerge in water. Density of water is 1 g/cc.

Grade:10

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Ajinkya Bhole,

Ans;- As it is clear from the problem that the weight of the block is (106-100) =6N=600 gm.

 it's vol is 10³=1000 cc

Hence the density is =600/1000=0.6 g/cc

Hence the fraction of the vol that will remain inside the water is =(0.6/1)=3/5.

So, the block will submerge upto a depth of 10×(3/5)=6 cm

 

  

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Ajinkya Bhole !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur